For a reaction, A + B ? Product; the rate law is given by, r = k [ A]1/2 [B]2. What is the order of the reaction?

Grade:8

1 Answers

Sunil Kumar FP
9 years ago
The order of a reaction with respect to a given substance is equal to its index or exponent to which its concentration term is raised
for in the rate law expression ,rate=k[A]x[B]y
the order of the reaction is x +y
according to the given problem
order=1/2+2=5/2